OpenStudy (anonymous):

pi is defined as the ratio of the circumference to the diameter of a circle, and it is an irrational number, right? It's value is evaluated to be 3.14...... and so on, and computers are still evaluating. My question is which circle's ratio of circumference to the diameter is being solved to obtain the value for pi? And what are those values of circumference and circle, which evaluates the value of pi?

OpenStudy (mr.math):

All circles!

OpenStudy (anonymous):

"Circumference of the circle" to the "diameter of the circle"

OpenStudy (anonymous):

@Mr. Math: All circles, right, so what are those values of circumference and diameter, which evaluates the value of pi?

OpenStudy (anonymous):

I mean, what number, do you divide with what number? as to get the answer 3.14......

OpenStudy (mr.math):

The ratio of the circumference of any circle to its diameter is pi.

OpenStudy (asnaseer):

you can read a little bit about how \(\pi\) was discovered here: http://www.gap-system.org/~history/HistTopics/Pi_through_the_ages.html

OpenStudy (mr.math):

You divide \(2\pi r\) by \(2r\) (i.e there's no restriction on the diameter or the circumference).

OpenStudy (anonymous):

Mr.Math is right - as usual :-P Just for the record, pi is irrational. You'll never get 3.14 exactly (unless by some coincidence sufficient rounding error lands you with this solution - but don't expect it!).

OpenStudy (turingtest):

So you got it, right? Circumference to diameter is\[\pi=\frac{C}{d}\]the circumfrance C is\[C=2\pi r\]and the diameter is\[d=2r\]so the ratio is\[\frac{C}{d}=\frac{2\pi r}{2r}=\pi\]The radius r cancels every time, regardless of the size of the circle...

OpenStudy (earthcitizen):

\[\tau= 2pi r\]

OpenStudy (earthcitizen):

\[\tau r=2\pi r\]

OpenStudy (anonymous):

I actually understand this, what I don't understand, is how computers and the scientists, achieved to find the value of pi, they how did they calculated?

OpenStudy (turingtest):

That is a much better question

OpenStudy (anonymous):

Read the article asnaseer linked, and otherwise search for "pi." I once saw a documentary on the subject - very fascinating - but I'm afraid I can't remember now.

OpenStudy (anonymous):

How do the computers evaluate the value? Some numbers have to be divided by some number to get the decimal places after 3.14? or even 3.14? What are they?

OpenStudy (turingtest):

I don't know exactly what king of programs are used, but I know there have been a variety of them, making use of thing like the way you can divide up a circle into triangles. Here are a list of many (doubtfully all) methods used to calculate pi: http://www.gap-system.org/~history/HistTopics/Pi_chronology.html but asnaseer's link is better for detail.

OpenStudy (asnaseer):

Another article on how to calculate \(\pi\): http://www.faqs.org/faqs/sci-math-faq/specialnumbers/computePi/

OpenStudy (turingtest):

...so in that link it says a computer used the "based on the Gauss-Legendre algorithm and Gauss arctangent relation" which I do not know the meaning of. But other methods used before may be more accessible.

OpenStudy (wasiqss):

i advise u not to go in such depth as it will screw u

OpenStudy (asnaseer):

@wasiqss - there is no harm in someone trying to find out more about the background of a particular subject or topic - it strengthens your knowledge by reinforcing your foundation.

OpenStudy (wasiqss):

lolz anaseer u lack xperience :P, i use to take keen interest bt i always ended up screwing, n the concepts i had also got vanished :P s

OpenStudy (asnaseer):

well - lets just say I will agree to disagree with you @wasiqss

OpenStudy (wasiqss):

well do wut u want who carex :P

OpenStudy (asnaseer):

@wasiqss - you need to have a more positive attitude towards things - it will help you calm down and understand things better :-)

OpenStudy (anonymous):

Thank you everyone for the help. It really helps. :)

OpenStudy (asnaseer):

yw

OpenStudy (wasiqss):

anaseer jus tel me wut grade u in?

OpenStudy (wasiqss):

n sorry if u did mind

OpenStudy (asnaseer):

@wasiqss - I'm currently working as a software engineer but I keep in touch with maths by /trying/ to teach others.

OpenStudy (wasiqss):

n im doin electronics engineerin :P

OpenStudy (asnaseer):

which year are you in?

OpenStudy (wasiqss):

1st

OpenStudy (asnaseer):

great! - I hope you enjoy your course